OpenStudy (anonymous):

How many possibilities are there if you are creating a 4-digit password using the digits 0 through 9? You may repeat numbers. A.40 B.6,561 C.5,040 D.10,000

OpenStudy (anonymous):

alot

OpenStudy (anonymous):

not an answer

OpenStudy (anonymous):

i know

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

don't think too hard for this one

OpenStudy (misty1212):

how many 4 digit numbers are there?

OpenStudy (anonymous):

B && C have 4 digit numbers

OpenStudy (anonymous):

my guess could be c but I'm not sure

OpenStudy (misty1212):

lol i was asking "how many four digit numbers are there" not 'how many answers have four digits in them?"

OpenStudy (misty1212):

ok lets go slow

OpenStudy (misty1212):

you have four digits to fill ten choices for each digit first digit 0 through 9 (ten choices) second digit another ten choices etc

OpenStudy (anonymous):

i have other questions do you know the aneswer

OpenStudy (misty1212):

by the counting principle, the number of choices all together is the product in other words, total number of choices is \[10\times 10\times 10\times 10=10^4\]

OpenStudy (misty1212):

which is another way of saying there are \(10,000\) four digit numbers

OpenStudy (anonymous):

i get it not thanks